# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ 4.1 Foundation Models Pretraining
We pretrain a large language model, denoted as $f_{\text{L}}^{\text{pre}}$ with parameters $\theta_{\text{L}}^{\text{pre}}$, using curated text corpora. The input data comprises omic entity descriptions $\mathcal{T}$, disease annotations $\mathcal{S}$, protein–protein interactions $\mathcal{E}^{(\text{PPI})}$, and drug–target relationships $\mathcal{E}^{(\text{DTI})}$. As illustrated in Figure 5, this pretraining phase equips the model with foundational knowledge of biomedical terminology and relational structure (see Appendix C.1), thereby enhancing its capacity for downstream reasoning in biomedical tasks. Then we will continue pretraining the language model $f_{\text{init}}$ parameterized by $\theta_{\text{init}}$, which are detailed in Appendix  C.2.
@@ 4.2 Model Training
The pipeline of model couples initial answering, a reinforcement-guided subgraph generator, and final answering. Given the structured prompt $P_{n}^{(\text{init})}$ from query $Q_{n}$, the pretrained model $f_{\text{init}}$ outputs $A_{n}^{(\text{init})}$ and a biomedical named entity recognition (NER) $\phi$ extracts entities $R_{n}^{(\text{init})}$ and maps them to proteins $\mathcal{V}_{n}^{(\text{init})}$, which will contribute to forming the start set $\mathcal{V}_{n}^{(\text{start})}$ depending on situation. Node features are embedded by pretrained graph foundation model to obtain $\mathcal{H}_{\text{in}}$ and formed $X_{n}^{(\text{cand})}$. Subgraph construction is framed as reinforcement learning problem consisting of four elements: state $\mathcal{G}_{n}^{(i)}=(\mathcal{V}_{n}^{(i)},\mathcal{E}_{n}^{(i)})$; action $\Delta_{n}^{(i)}=(v_{\text{src}}^{i},v_{\text{tgt}}^{i})$ adds a single edge under feasibility masks; policy uses a message propagation (MSG) module to produce $X_{n}^{(i)}=\pi_{\text{MSG}}(\mathcal{G}_{n}^{(i)},X_{n}^{(\text{cand})})$ and two masked probability function $\pi_{\text{SRC}},\pi_{\text{TGT}}$ to sample $v_{\text{src}}^{i}$ and $v_{\text{tgt}}^{i}$; reward combines feedback from a pretrained classifier $g(\cdot)$, a rollout averaging $L$ simulated continuations, and a rule-based term $\mathcal{R}_{\text{rule}}$ that penalizes schema violations. We accept an action e only when $\mathcal{R}^{(i)}_{\text{total}}>0$ and update the generator with reward-weighted cross-entropy with early stopping; the best subgraph $\mathcal{G}_{n}^{\dagger}$ is retained in $\Omega$ stochastic run. For final answering, $\mathcal{G}_{n}^{\dagger}$ is verbalized in expert mode  (Fatemi et al., 2023) and appended to $Q_{n}$ to form $P_{n}^{(\text{final})}$; the model $f_{\text{final}}$ is finetuned with token-level cross-entropy against $A_{n}$. What follows is a detailed exposition of the model design and training framework.
@@ 5 Experiments
We construct the Target-QA dataset over multiple TCGA cancer types using DepMap, integrating multi-omic features (epigenomic, genomic, transcriptomic, proteomic) and metadata from cancer cell lines. Each QA pair consists of an input query, including multi-omic and cell line information, and an output answer comprising the top-$\gamma$ ($\gamma=100$) CRISPR-prioritized targets. The final dataset contains 363 QA pairs from cancerous cell lines after integration and preprocessing. We use an 80/20 train-test split and repeat experiments across four randomized seeds to ensure stability and generalization. For foundation model pretraining, we collect dataset of 336 samples with multi-omic features from unannotated samples, including both disease and control groups (297 cancerous, 39 non-cancerous), and apply stratified sampling to address class imbalance. Full data processing and cohort composition are described in Appendix  B).
@@
We initialize the language model with LLaMA3-8B-Instruct (AI@Meta, 2024), pretrained on biomedical terminology and curated textual descriptions involving protein–protein and disease–protein relationships from BioMedGraphica to enhance domain-specific vocabulary and biological context understanding (see Figure 5). For graph encoding, we use BioBERT-v1.1 (Lee et al., 2020) for text embeddings and Graph Attention Networks (GAT) (Veličković et al., 2017) to learn topological features from protein interaction graphs, incorporating random edge masking to improve robustness (see Figure 2 global message propagation). The pretrained GNN achieves 64.4% AUC in edge prediction, and 99.46% / 96.15% accuracy on disease type classification (train/test). Named entities are extracted using GPT-4o-mini via ChatGPT API (OpenAI, 2023).GALAX is trained using the Adam optimizer on two NVIDIA H100 GPUs (80GB). We set the number of top omic features per modality to K=10K{=}10, the maximum subgraph rollout depth to L=5L{=}5, and the number of candidate starting nodes $\eta{=}20$. The reward formulation includes both rollout- and rule-based components, each weighted equally with $\lambda{=}1$ and $\lambda_{\text{rule}}{=}1$. The reasoning task is formulated as a binary classification problem ($|\mathcal{O}|{=}2$), using a 1-hop (h=1h{=}1) protein neighborhood from the disease-annotated subgraph. Model outputs are evaluated using precision, recall, F1-score, Jaccard similarity, Hit@5 and Hit@10, by comparing predicted target sets $\hat{R}_{n}$ against reference labels $R_{n}$. Additional details are provided in the Appendix D.2.
@@
L3 + Omics + KG $0125\scriptstyle\pm 0.0032 0.0$$029\scriptstyle\pm 0.0003 0.0$$014\scriptstyle\pm 0.0025 0.0$$010\scriptstyle\pm 0.0017 0.0$$073\scriptstyle\pm 0.0068 0.0$$0033\scriptstyle\pm 0.0029$
@@
L3-FT(Med) + Omics + KG $0158\scriptstyle\pm 0.0030 0.0$$058\scriptstyle\pm 0.0011 0.0$$081\scriptstyle\pm 0.0071 0.0$$024\scriptstyle\pm 0.0017 0.0$$149\scriptstyle\pm 0.0057 0.0$$0050\scriptstyle\pm 0.0000$
@@
GALAX $\mathbf{0.5472}\scriptstyle\pm\mathbf{0.0053}$ $\mathbf{0.5332}\scriptstyle\pm\mathbf{0.0031}$ $\mathbf{0.5345}\scriptstyle\pm\mathbf{0.0185}$ $\mathbf{0.5157}\scriptstyle\pm\mathbf{0.0043}$ $\mathbf{0.5608}\scriptstyle\pm\mathbf{0.0031}$ $\mathbf{0.5533}\scriptstyle\pm\mathbf{0.0033}$
@@
L3-FT(QA) + Omics $0.8693\scriptstyle\pm 0.0157$ $0.8889\scriptstyle\pm 0.0168$ $0.8667\scriptstyle\pm 0.0218$ $0.8476\scriptstyle\pm 0.0165$ $0.8389\scriptstyle\pm 0.0096$ $0.8889\scriptstyle\pm 0.0509$
@@
GALAX $\mathbf{0.8815}\scriptstyle\pm\mathbf{0.0033}$ $\mathbf{0.9249}\scriptstyle\pm\mathbf{0.0048}$ $\mathbf{0.8810}\scriptstyle\pm\mathbf{0.0082}$ $\mathbf{0.9238}\scriptstyle\pm\mathbf{0.0436}$ ${0.8500}\scriptstyle\pm{0.0441}$ $\mathbf{0.8889}\scriptstyle\pm\mathbf{0.0839}$
Baselines and Ablation Study
Traditional method, M2T (Multiomic2Target(Deng et al., 2024)), serves as a baseline that uses only multi-omic features without graph or language modeling and performs poorly across all metrics. And we perform ablation studies to isolate the contribution of core components of language and graph modules in GALAX. On the language axis, a non–task-tuned LLaMA3 (L3+Omics) is weak; domain-adaptive finetuning on biomedical text (L3-FT(Med)+Omics) yields modest gains; task-adaptive finetuning on Target-QA (L3-FT(QA)+Omics) produces the step-change. On the graph axis, GAT incorporates graph foundation model and trained to predict the CRISPR knockout effects, shows limited improvements. Furthermore, integrating graph modules into language models comes with different outputs. Adding a static KG to each language foundation models ([L3 / L3-FT(Med) / L3-FT(QA)]+Omics+KG) barely improve or even decrease model performances, and graph retrieval with pretrained GAT (G-Retriever+pre-GAT) outperform some task-adaptive finetuned language models but not reliably due to the difficulty of extracting relevant subgraphs from millions of nodes/edges. Reinforcement-guided subgraph construction on top of QA-tuned language (GALAX: L3-FT(QA)+KG+RL) delivers consistent, cross-dataset gains of roughly 2%-5% on each metric, indicating that a reinforcement-guided subgraph generator outperforms other graph augmented models across all datasets by enabling process-level reasoning over biologically plausible subgraphs, and that QA finetuning provides the dominant improvement (shown in Tables 12). Full baseline and ablation results are provided in Appendix  D.3.
Computational Complexity
@@
We denote the language model complexity as $O(\kappa)$. The retrieved KG subgraph contains $M$ nodes, and $\varepsilon$ denotes graph embedding cost. Augmented with KG retrieval, L3-FT(QA)+Omics+KG adds $O(M^{2})$ per query and results in $O(\kappa+M^{2}\varepsilon)$ for training and inference. G-retriever+preGAT embed $M$ nodes, requiring $O(M\varepsilon+M^{2}\varepsilon)$, thus $O(\kappa+M\varepsilon+M^{2}\varepsilon)$ for training and inference. GALAX augments language model with reinforcement-guided subgraph construction, whose embedding costs $O(M\varepsilon+M^{2}\varepsilon)$. The LLM demands an $O(\kappa)$ initialization for the top $\eta$ candidates and the same for final answer generation; since $\eta\ll M$, graph embedding dominates the reward term in the RL cost. Consequently, the overall training and inference complexity is $O(\kappa+M\varepsilon+M^{2}\varepsilon)$ (see Table 3).

### v2

@@ 4.1 Foundation Models Pretraining
We pretrain a large language model, denoted as $f_{\text{L}}^{\text{pre}}$ with parameters $\theta_{\text{L}}^{\text{pre}}$, using curated text corpora. The input data comprises omic entity descriptions $\mathcal{T}$, disease annotations $\mathcal{S}$, protein–protein interactions $\mathcal{E}^{(\text{PPI})}$, and disease–target relationships $\mathcal{E}^{(\text{DTI})}$. As illustrated in Figure 5, this pretraining phase equips the model with foundational knowledge of biomedical terminology and relational structure (see Appendix C.1), thereby enhancing its capacity for downstream reasoning in biomedical tasks. Then we will continue pretraining the language model $f_{\text{init}}$ parameterized by $\theta_{\text{init}}$, which are detailed in Appendix  C.2.
@@ 4.2 Model Training
The pipeline of model couples initial answering, a reinforcement-guided subgraph generator, and final answering. Given the structured prompt $P_{n}^{(\text{init})}$ from query $Q_{n}$, the pretrained model $f_{\text{init}}$ outputs $A_{n}^{(\text{init})}$ and a biomedical named entity recognition (NER) $\phi$ extracts entities $R_{n}^{(\text{init})}$ and maps them to proteins $\mathcal{V}_{n}^{(\text{init})}$, which will contribute to forming the start set $\mathcal{V}_{n}^{(\text{start})}$ depending on situation. Node features are embedded by pretrained graph foundation model to obtain $\mathcal{H}_{\text{in}}$ and formed $X_{n}^{(\text{cand})}$. Subgraph construction is framed as reinforcement learning problem consisting of four elements: state $\mathcal{G}_{n}^{(i)}=(\mathcal{V}_{n}^{(i)},\mathcal{E}_{n}^{(i)})$; action $\Delta_{n}^{(i)}=(v_{\text{src}}^{i},v_{\text{tgt}}^{i})$ adds a single edge under feasibility masks; policy uses a message propagation (MSG) module to produce $X_{n}^{(i)}=\pi_{\text{MSG}}(\mathcal{G}_{n}^{(i)},X_{n}^{(\text{cand})})$ and two masked probability function $\pi_{\text{SRC}},\pi_{\text{TGT}}$ to sample $v_{\text{src}}^{i}$ and $v_{\text{tgt}}^{i}$; reward combines feedback from a pretrained classifier $g(\cdot)$, a rollout averaging $L$ simulated continuations, and a rule-based term $\mathcal{R}_{\text{rule}}$ that penalizes schema violations. We accept an action e only when $\mathcal{R}^{(i)}_{\text{total}}>0$ and update the generator with reward-weighted cross-entropy with early stopping; the best subgraph $\mathcal{G}_{n}^{\dagger}$ is retained in $\Omega$ stochastic run. For final answering, $\mathcal{G}_{n}^{\dagger}$ is verbalized in expert mode  (fatemi2023talk) and appended to $Q_{n}$ to form $P_{n}^{(\text{final})}$; the model $f_{\text{final}}$ is finetuned with token-level cross-entropy against $A_{n}$. What follows is a detailed exposition of the model design and training framework.
@@ 5 Experiments
We construct the Target-QA dataset over multiple TCGA cancer types using DepMap, integrating multi-omic features (epigenomic, genomic, transcriptomic, proteomic) and metadata from cancer cell lines. Each QA pair consists of an input query, including multi-omic and cell line information, and an output answer comprising the top-$\gamma$ ($\gamma=100$) CRISPR-prioritized targets. The final dataset contains 363 QA pairs from cancerous cell lines after integration and preprocessing. We use an 80/20 train-test split and repeat experiments across four randomized seeds to ensure stability and generalization. For foundation model pretraining, we collect dataset of 336 samples with multi-omic features from unannotated samples, including both disease and control groups (297 cancerous, 39 non-cancerous), and apply stratified sampling to address class imbalance. Full data processing and cohort composition are described in Appendix  B.
@@
We initialize the language model with LLaMA3-8B-Instruct (llama3modelcard), pretrained on biomedical terminology and curated textual descriptions involving protein–protein and disease–protein relationships from BioMedGraphica to enhance domain-specific vocabulary and biological context understanding (see Figure 5). For graph encoding, we use BioBERT-v1.1 (lee2020biobert) for text embeddings and Graph Attention Networks (GAT) (velivckovic2017graph) to learn topological features from protein interaction graphs, incorporating random edge masking to improve robustness (see Figure 2 global message propagation). The pretrained GNN achieves 64.4% AUC in edge prediction, and 99.46% / 96.15% accuracy on disease type classification (train/test). Named entities are extracted using GPT-4o-mini via ChatGPT API (chatgpt2023).GALAX is trained using the Adam optimizer on two NVIDIA H100 GPUs (80GB). We set the number of top omic features per modality to K=10K{=}10, the maximum subgraph rollout depth to L=5L{=}5, and the number of candidate starting nodes $\eta{=}20$. The reward formulation includes both rollout- and rule-based components, each weighted equally with $\lambda{=}1$ and $\lambda_{\text{rule}}{=}1$. The reasoning task is formulated as a binary classification problem ($|\mathcal{O}|{=}2$), using a 1-hop (h=1h{=}1) protein neighborhood from the disease-annotated subgraph. Model outputs are evaluated using precision, recall, F1-score, Jaccard similarity, Hit@5 and Hit@10, by comparing predicted target sets $\hat{R}_{n}$ against reference labels $R_{n}$. Additional details are provided in the Appendix D.2.
@@
L3 + Omics + KG $0125\scriptstyle\pm 0.0032 0.0$$029\scriptstyle\pm 0.0003 0.0$$014\scriptstyle\pm 0.0025 0.0$$010\scriptstyle\pm 0.0016 0.0$$073\scriptstyle\pm 0.0068 0.0$$0033\scriptstyle\pm 0.0029$
@@
L3-FT(Med) + Omics + KG $0158\scriptstyle\pm 0.0030 0.0$$058\scriptstyle\pm 0.0011 0.0$$081\scriptstyle\pm 0.0071 0.0$$024\scriptstyle\pm 0.0016 0.0$$149\scriptstyle\pm 0.0057 0.0$$0050\scriptstyle\pm 0.0000$
@@
RoG $0.5248\scriptstyle\pm 0.0134$ $0.4726\scriptstyle\pm 0.0445$ $0.5213\scriptstyle\pm 0.0227$ $0.4562\scriptstyle\pm 0.0848$ $0.4791\scriptstyle\pm 0.0575$ $0.4311\scriptstyle\pm 0.0721$
SubgraphRAG $0.5280\scriptstyle\pm 0.0044$ $0.4617\scriptstyle\pm 0.0027$ $0.5123\scriptstyle\pm 0.0105$ $0.4448\scriptstyle\pm 0.0386$ $0.4708\scriptstyle\pm 0.0317$ $0.3917\scriptstyle\pm 0.0376$
GNN-RAG $0.5258\scriptstyle\pm 0.0126$ $0.4735\scriptstyle\pm 0.0190$ $334\scriptstyle\pm 0.0225 0.505$$052\scriptstyle\pm 0.0170$ $0.4787\scriptstyle\pm 0.0453$ $0.4389\scriptstyle\pm 0.0584$
GALAX $\mathbf{0.5472}\scriptstyle\pm\mathbf{0.0053}$ $0.5332\scriptstyle\pm 0.0031$ $0.5345\scriptstyle\pm 0.0185$ $0.5157\scriptstyle\pm 0.0043$ $\mathbf{0.5608}\scriptstyle\pm\mathbf{0.0031}$ $\mathbf{0.5533}\scriptstyle\pm\mathbf{0.0033}$
GALAX (Qwen2.5-7B) $0.5445\scriptstyle\pm 0.0114$ $\mathbf{0.5405}\scriptstyle\pm\mathbf{0.0101}$ $\mathbf{0.5475}\scriptstyle\pm\mathbf{0.0019}$ $\mathbf{0.5462}\scriptstyle\pm\mathbf{0.0111}$ $0.5171\scriptstyle\pm 0.0474$ $0.5206\scriptstyle\pm 0.0419$
@@
L3-FT(QA) + Omics $0.8693\scriptstyle\pm 0.0157$ $0.8889\scriptstyle\pm 0.0168$ $0.8667\scriptstyle\pm 0.0218$ $0.8476\scriptstyle\pm 0.0165$ $0.8389\scriptstyle\pm 0.0096$ $\mathbf{0.8889}\scriptstyle\pm\mathbf{0.0509}$
@@
RoG $0.8450\scriptstyle\pm 0.0350$ $0.8593\scriptstyle\pm 0.0318$ $0.8238\scriptstyle\pm 0.0218$ $0.8095\scriptstyle\pm 0.0436$ $11\scriptstyle\pm 0.1110 0.7667±0$$67\scriptstyle\pm 0.0577$
SubgraphRAG $0.8476\scriptstyle\pm 0.0167$ $0.8624\scriptstyle\pm 0.0120$ $0.8238\scriptstyle\pm 0.0082$ $0.8190\scriptstyle\pm 0.0165$ $0.7333\scriptstyle\pm 0.1014$ $0.7556\scriptstyle\pm 0.0839$
GNN-RAG $0.8323\scriptstyle\pm 0.0205$ $0.8656\scriptstyle\pm 0.0302$ $0.7571\scriptstyle\pm 0.0623$ $0.7905\scriptstyle\pm 0.0719$ $0.8222\scriptstyle\pm 0.0674$ $0.8444\scriptstyle\pm 0.0385$
GALAX $0.8815\scriptstyle\pm 0.0033$ $\mathbf{0.9249}\scriptstyle\pm\mathbf{0.0048}$ $\mathbf{0.8810}\scriptstyle\pm\mathbf{0.0082}$ $\mathbf{0.9238}\scriptstyle\pm\mathbf{0.0436}$ $0.8500\scriptstyle\pm 0.0441$ $\mathbf{0.8889}\scriptstyle\pm\mathbf{0.0839}$
GALAX (Qwen2.5-7B) $\mathbf{0.8841}\scriptstyle\pm\mathbf{0.0126}$ $0.9079\scriptstyle\pm 0.0084$ $0.8667\scriptstyle\pm 0.0082$ $0.9048\scriptstyle\pm 0.0165$ $0.8000\scriptstyle\pm 0.0764$ $0.8556\scriptstyle\pm 0.0385$
Baseline Models
Traditional method, M2T (Multiomic2Target(deng2024multiomics2targets)), serves as a baseline that uses only multi-omic features without graph or language modeling and performs poorly across all metrics. And we perform ablation studies to isolate the contribution of core components of language and graph modules in GALAX. On the language axis, a non–task-tuned LLaMA3 (L3+Omics) is weak; domain-adaptive finetuning on biomedical text (L3-FT(Med)+Omics) yields modest gains; task-adaptive finetuning on Target-QA (L3-FT(QA)+Omics) produces the step-change. On the graph axis, GAT incorporates graph foundation model and trained to predict the CRISPR knockout effects, shows limited improvements. Furthermore, integrating graph modules into language models comes with different outputs. Adding a static KG to each language foundation models ([L3 / L3-FT(Med) / L3-FT(QA)]+Omics+KG) barely improve or even decrease model performances, and graph retrieval with pretrained GAT (G-Retriever+pre-GAT) outperform some task-adaptive finetuned language models but not reliably due to the difficulty of extracting relevant subgraphs from millions of nodes/edges. RoG, SubgraphRAG (li2024simple), and GNN-RAG (mavromatis2025gnn) augment the language model by retrieving optimal paths, achieving moderate improvements over L3-FT(QA)+Omics+KG. Reinforcement-guided subgraph construction on top of QA-tuned language (GALAX: L3-FT(QA)+Omics+KG+RL) delivers consistent, cross-dataset gains of roughly 2%-5% on each metric, indicating that a reinforcement-guided subgraph generator outperforms other graph augmented models across all datasets by enabling process-level reasoning over biologically plausible subgraphs (shown in Tables 12). Details of baseline models are provided in Appendix  D.3.
@@
RoG $O(\kappa+M\varepsilon+M^{2}\varepsilon)$
GNN-RAG $O(\kappa+M\varepsilon+M^{2}\varepsilon)$
@@
Computational Complexity
We denote the language model complexity by $O(\kappa)$. The retrieved KG subgraph includes $M$ nodes, and $\varepsilon$ denotes the graph embedding cost. When augmented with KG retrieval, L3-FT(QA)+Omics+KG introduces an additional $O(M^{2})$ per query, yielding $O(\kappa+M^{2})$ for both training and inference. G-retriever+preGAT embeds all $M$ nodes and thus incurs $O(M\varepsilon+M^{2}\varepsilon)$. RoG and GNN-RAG follow the same cost at retrieval, since they embed entities and relations, requiring $O(M\varepsilon+M^{2}\varepsilon)$. SubgraphRAG reduces this by retrieving only relation triplets, which costs $O(M^{2}\varepsilon)$ at retrieval. And G-retreiver. RoG, GNN-RAG and SubgraphRAG all requires an $O(\kappa)$ at both training and inference for language models. GALAX augments the language model with reinforcement-guided subgraph construction, with embedding cost $O(M\varepsilon+M^{2}\varepsilon)$. The model requires an $O(\kappa)$ forward pass to initialize the top $\eta$ candidates and another for final answer generation; since $\eta\ll M$, the graph-embedding term dominates the RL reward cost. Overall, the training and inference complexity of GALAX is therefore $O(\kappa+M\varepsilon+M^{2}\varepsilon)$ (see Table 3).